### v1

# Multi-Agent Q-Learning Dynamics in Random Networks: Convergence due to Exploration and Sparsity

Abstract
Beyond specific settings, many multi-agent learning algorithms fail to converge to an equilibrium solution, and instead display complex, non-stationary behaviours such as recurrent or chaotic orbits. In fact, recent literature suggests that such complex behaviours are likely to occur when the number of agents increases. In this paper, we study Q-learning dynamics in network polymatrix games where the network structure is drawn from classical random graph models. In particular, we focus on the Erdős-Rényi model, a well-studied model for social networks, and the Stochastic Block model, which generalizes the above by accounting for community structures within the network. In each setting, we establish sufficient conditions under which the agents’ joint strategies converge to a unique equilibrium. We investigate how this condition depends on the exploration rates, payoff matrices and, crucially, the sparsity of the network. Finally, we validate our theoretical findings through numerical simulations and demonstrate that convergence can be reliably achieved in many-agent systems, provided network sparsity is controlled.
1 Introduction
The development of algorithms for multi-agent learning has produced a number of successes in recent years, solving challenging problems in load-balancing (Krichene et al., 2014; Southwell et al., 2012; Moghaddam & Kebriaei, 2024), energy management (Paccagnan et al., 2018; Maddouri et al., 2018) and game playing (Moravčík et al., 2017; Brown & Sandholm, 2019; Samvelyan et al., 2019; Perolat et al., 2022). Also thanks to these successes, the game-theoretical foundations of learning in the face of many agents remain a thriving area of research. An important question is whether learning algorithms are guaranteed to converge to an equilibrium solution, often the Nash Equilibrium. However, such guarantees exist only for specialized settings, such as two-player games (Metrick & Polak, 1994; Kianercy & Galstyan, 2012), zero-sum network games (Ewerhart & Valkanova, 2020; Leonardos et al., 2024) or potential games (Krichene et al., 2015; Harris, 1998). Often, such assumptions are restrictive and do not hold in practice. As the development of learning in the presence of many agents progresses, it therefore becomes necessary to understand whether convergence guarantees hold outside these restrictive assumptions.
Unfortunately, it is precisely at this point where results on non-convergence appear. A strong example is Milionis et al. (2023), which introduced a game in which no independent learning dynamics converges to a Nash Equilibrium. Further studies have shown that chaotic dynamics may occur even in two-player finite-action games (Sato et al., 2002; Galla & Farmer, 2013). Crucially, Sanders et al. (2018) found that, as the number of players grows, non-convergent behaviour becomes the norm.
Whilst at first glance these results seem to suggest an insurmountable barrier towards strong guarantees of convergence, a key missing factor is an in-depth analysis of the interactions between agents. In particular, both Sanders et al. (2018) and Hussain et al. (2023) assume that the payoff to any given agent is dependent on all other agents in the environment. Yet, in practice, this assumption rarely holds. Rather, agents often interact through an underlying network that may represent social connections, communication constraints, or spatial proximity. In this work, we address this concern by studying learning in network games (Kearns et al., 2001). In particular, we are motivated by the following research question:
How does network structure affect the convergence of multi-agent learning as the number of agents increases?
Answering this question represents a key step towards guaranteeing the feasibility of learning with many agents, so long as the network structure can be controlled. Indeed, a number of works have examined learning in network games, uncovering the relationship between the network structure and properties of the equilibrium as well as designing algorithms that converge to an equilibrium (Melo, 2018; Parise & Ozdaglar, 2019; Melo, 2021; Shokri & Kebriaei, 2020). However, these algorithms require full knowledge of the agents’ payoff functions or evaluations of its gradient, which may not be available in practice. Our goal instead is to consider a widely applied reinforcement learning algorithm – Q-Learning – which requires only evaluations of the payoff functions, and study how the parameters of the algorithm and the structure of the underlying network can be leveraged to yield convergence guarantees, as well as how such guarantees scale with the number of agents.
Model.
We study agents who update their strategies via the Q-Learning dynamics (Sato & Crutchfield, 2003; Tuyls et al., 2006), a continuous-time counterpart to the well-studied Q-Learning algorithm (Watkins & Dayan, 1992; Sutton & Barto, 2018). We focus our study on network polymatrix games (Janovskaja, 1968) in which agents select from a finite set of actions and interact with their neighbours on an underlying network. We study networks which are drawn from random graph models, where the sparsity of the network is parameterised by the distribution from which the network is drawn: the Erdős-Rényi model (Erdös & Rényi, 2006; Newman, 2010) and the Stochastic Block Model (Holland et al., 1983). The Erdős-Rényi model is widely used to study social networks and natural phenomena such as information diffusion Li et al. (2019). The Stochastic Block Model is a natural extension of Erdős-Rényi to model community structures which naturally arise in distributed systems (Yun & Proutière, 2019).
Contributions.
Our main contribution is to provide a precise characterisation of the convergence properties of Q-Learning, depending on the exploration rates of the agents, the payoffs in the game, and the sparsity of the network. Specifically, our results provide sufficient conditions for the Q-learning dynamics to converge to an equilibrium with high probability as the number of agents increases, given that the network is drawn from one of the two above-mentioned models. We show that, in sparsely connected networks, Q-learning converges even with low exploration rates as the number of agents increases, whereas, in agreement with previous results, learning in densely connected networks may not converge. In doing so, we show an explicit relationship between the convergence the Q-Learning Dynamics and the sparsity of the network, thus ensuring the feasibility of learning in many-agent games, so long as the network sparsity is controlled. To the best of our knowledge, ours is the first work to study the asymptotic behaviour of learning dynamics in the context of network polymatrix game with a random graph model, and to derive the relation between the sparsity of the network.
As part of our results, we show that our sufficient conditions also guarantee the uniqueness of the equilibrium. In this way, we show not only that Q-Learning converges, but also that it does so to a unique solution, for all initial conditions. We finally adapt our results to determine a condition that guarantees the uniqueness of the Nash Equilibrium.
Related Work
Our work is focused on independent online learning in network polymatrix games. Network games are well-studied within the setting of zero-sum networks (Cai et al., 2016), which model strictly competitive systems. In such cases, it is known that the continuous-time counterparts of popular algorithms such as Fictitious Play (Ewerhart & Valkanova, 2020) and Q-Learning (Leonardos et al., 2024) converge to an equilibrium. By contrast it was shown in Bailey & Piliouras (2019) and Abernethy et al. (2021) that the Replicator Dynamics (Maynard Smith, 1974), a continuous-time model of the Multiplicative Weights Update algorithm (Arora et al., 2012) does not converge to the NE, but rather cycles at a fixed distance away from it.
Network Games have also been studied to understand the properties, in particular the uniqueness, of the equilibrium (Bramoullé et al., 2014; Parise & Ozdaglar, 2019; Melo, 2021). In many of these cases, the literature appeals to the study of monotone games (Paccagnan et al., 2018) which subsumes zero-sum network games (Akin & Losert, 1984). The formalism of monotone games has been applied to design algorithms that provably converge to Nash Equilibria. However, many of these require that agents have full access to their payoff function (Parise & Ozdaglar, 2019) or its evaluations of its gradient (Mertikopoulos & Zhou, 2019). Monotone games also share strong links with the idea of payoff perturbations (Facchinei & Pang, 2004) in which a strongly convex penalty is imposed to agents’ payoff functions to stabilise learning (Abe et al., 2024; Sokota et al., 2023; Liu et al., 2023).
Our work departs from the above by considering Q-Learning Dynamics, a foundational exploration and exploitation model central to reinforcement learning theory (Albrecht et al., 2024a; Tuyls, 2023). Q-Learning Dynamics is also known to be related to the replicator dynamics (Bloembergen et al., 2015) and to Follow-the-Regularised-Leader (Abe et al., 2022). Outside of specialised classes of games, Q-Learning Dynamics has been shown to exhibit chaotic orbits (Sato et al., 2002; Galla & Farmer, 2013), a phenomenon which becomes more prevalent as the number of players increases (Sanders et al., 2018). Similar to our work, Hussain et al. (2024) consider the convergence of Q-Learning in network polymatrix games. However, their work considered deterministic graphs with specific structures. We instead consider Q-Learning in a stochastic setting, where the stochasticity comes from the random network. In doing so, we derive a direct relationship between the sparsity of edge connections in the network and the convergence of Q-Learning Dynamics.
To achieve our result, we appeal to the framework of random networks, which are often used to model social and community structures. Many works in multi-agent learning, and online optimisations have considered this setting, for example to determine the existence of pure Nash equilibria (Daskalakis et al., 2011), to study the performance of distributed online algorithms (Lei et al., 2020), or to examine the emergence of cooperative behaviours (Dall’Asta et al., 2012). An important step in the study of network games was the introduction of Graphon games in Parise & Ozdaglar (2023), which are the limit of network games as the agents tend to infinity. The authors show that graphons generalise Erdős-Rényi and Stochastic Block Models. However, the restriction to the limit of infinite players is not suitable for our purposes, as it does not allow us to examine how our convergence guarantees are influenced by the number of agents.
2 Preliminaries
We begin in Sec. 2.1 by describing network games. In Sec.  2.2, we describe the equilibrium concepts which are of particular interest in multi-agent learning. We then describe Q-Learning in Sec. 2.3.
2.1 Game Model
In this work, we consider network polymatrix games Janovskaja (1968); Cai et al. (2016), which are defined as tuples $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A}_{k})_{k\in\mathcal{N}},(A^{kl},A^{lk})_{(k,l)\in\mathcal{E}})$, in which $\mathcal{N}=\{1,2,\ldots,N\}$ denotes a set of $N$ agents, indexed by $k$. The interactions between agents is modelled by a set $\mathcal{E}\subset\mathcal{N}\times\mathcal{N}$ of edges that defines an undirected network. An alternative formulation of the underlying network is through an adjacency matrix $G\in\mathbb{R}^{N\times N}$ in which $[G]_{kl}=1$ if $(k,l)\in\mathcal{E}$, and $[G]_{kl}=0$ otherwise.
At each round, each agent $k$ selects an action $i\in\mathcal{A}_{k}$, where $\mathcal{A}_{k}$ is a finite set of $n_{k}$ actions. We denote the strategy of an agent as the probability distribution over their actions. Next, we define the joint strategy across all agents as the concatenation of all individual strategies $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ and apply the shorthand $\mathbf{x}_{-k}$ to denote the strategies of all agents other than $k$.
The goal of each agent is to maximise a utility function $u_{k}$. In a network polymatrix game $\mathcal{G}$, each edge is associated with the payoff matrices $(A^{kl},A^{lk})$ so that the payoff for each agent $k$ takes the form
$u_{k}(\mathbf{x}_{k},\mathbf{x}_{-k})=\sum_{l:(k,l)\in\mathcal{E}}\mathbf{x}_{k}^{\top}A^{kl}\mathbf{x}_{l}.$
2.2 Solution Concepts
In this work, we are concerned with two widely-studied solution concepts: the Nash Equilibrium (Nash, 1950) and the Quantal Response Equilibrium (McKelvey & Palfrey, 1995). To define these concepts we first define the reward to agent $k\in\mathcal{N}$ for playing action $i\in\mathcal{A}_{k}$ as $r_{ki}(\mathbf{x}_{-k})=\sum_{l:(k,l)\in\mathcal{E}}\sum_{j\in\mathcal{A}_{l}}[A^{kl}]_{ij}x_{lj}$.
Definition 1 (Nash Equilibrium).
A joint strategy $\mathbf{x}^{*}\in\Delta(\mathcal{A})$ is a Nash Equilibrium if, for all $k\in\mathcal{N}$,
$\mathbf{x}_{k}^{*}\in\operatorname*{arg\,max}_{\mathbf{x}_{k}\in\Delta(\mathcal{A}_{k})}u_{k}(\mathbf{x}_{k},\mathbf{x}_{-k}^{*}).$
Definition 2 (Quantal Response Equilibrium (QRE)).
Let $T_{1},\ldots,T_{N}\geq 0$. Then, a joint strategy $\mathbf{x}^{*}\in\Delta(\mathcal{A})$ is a Quantal Response Equilibrium if, for all $k\in\mathcal{N}$,
$\mathbf{x}_{k}^{*}=\frac{\exp\left(r_{ki}(\mathbf{x}_{-k})/T_{k}\right)}{\sum_{j\in\mathcal{A}_{k}}\exp\left(r_{kj}(\mathbf{x}_{-k})/T_{k}\right)}.$
The QRE is a natural extension of the Nash Equilibrium that allows for agents to play suboptimal actions with non-zero probability. This is crucial in online learning, where agents must explore their strategy space. The parameter $T_{k}$ is therefore known as the exploration rate of agent $k$. Notice that, by taking the limit $T_{k}\rightarrow 0$ for all $k$, the QRE converges to the Nash Equilibrium (McKelvey & Palfrey, 1995). On the other hand, as $T_{k}\rightarrow\infty$ for all $k$, the QRE converges to the uniform distribution, in which payoffs are ignored and agents play their actions at random. For this reason, it is preferred that the agents learn QRE at lower values of $T_{k}$, so that the learned solution approaches the Nash Equilibrium (Hussain et al., 2024).
2.3 Q-Learning Dynamics
We now describe the Q-Learning algorithm (Watkins & Dayan, 1992) which aims to learn an optimal action-value function $Q:\mathcal{A}\rightarrow\mathbb{R}$ that captures the expected reward of taking a given action. We consider the multi-agent extension of Q-Learning (Schwartz, 2014; Albrecht et al., 2024b) in which each agent $k$ maintains an individual Q-value estimate $Q_{k}:\mathcal{A}_{k}\rightarrow\mathbb{R}$. These are updated at each round $t$ via the update
$Q_{ki}(t+1)=(1-\alpha_{k})Q_{ki}(t)+\alpha_{k}r_{ki}(\mathbf{x}_{-k}(t))$
where $\alpha_{k}\in(0,1)$ is the learning rate of agent $k$.
In effect, $Q_{ki}$ gives a discounted history of the rewards received when action $i$ is played, with $1-\alpha_{k}$ as the discount factor. It should also be noted that Q-values are updated by the rewards $r_{ki}(\mathbf{x}_{-k})$ that depend on the strategies of all other agents at time $t$. So, the reward for a single action can vary between rounds. It is this non-stationarity that often leads to chaotic dynamics in multi-agent learning (Sato et al., 2002).
Given the Q-values, each agent plays their actions according to the Boltzmann distribution (Sutton & Barto, 2018) given by
$x_{ki}(t+1)=\frac{\exp\left(Q_{ki}(t+1)/T_{k}\right)}{\sum_{j\in\mathcal{A}_{k}}\exp\left(Q_{kj}(t+1)/T_{k}\right)}$ (1)
where $T_{k}\in(0,\infty)$ is the exploration rate of agent $k$.
Tuyls et al. (2006) and Sato & Crutchfield (2003) showed that a continuous-time approximation of the Q-Learning algorithm is given by a variation of the replicator dynamics (Maynard Smith, 1974; Hofbauer & Sigmund, 2003) that is called the Q-Learning Dynamics
$\frac{\dot{x}_{ki}}{x_{ki}}=r_{ki}\left(\mathbf{x}_{-k}\right)-\langle\mathbf{x}_{k},r_{k}(\mathbf{x})\rangle+T_{k}\sum_{j\in\mathcal{A}_{k}}x_{kj}\ln\frac{x_{kj}}{x_{ki}},$ (QLD)
[…]
Additional Notation.
Given a square matrix $A\in\mathbb{R}^{N\times N}$, we denote its spectral norm as $\lVert A\rVert_{2}=\sup_{x\in\mathbb{R}^{N}\,:\,\lVert\mathbf{x}\rVert_{2}=1}\lVert Ax\rVert_{2}$. If $A$ is symmetric, we denote its spectral radius as $\rho(G)=\max_{i=1,\ldots,N}|\lambda_{i}|$, where $\lambda_{1},\ldots,\lambda_{N}$ denote the eigenvalues of $A$.
3 Convergence of Q-Learning in Graphs
In this section, we analyse the relation between network topology and the convergence of (QLD). In particular, in Lemma 1 we determine a sufficient condition on the exploration rates $T_{k}$ such that (QLD) converges to the unique QRE with high probability as $N\rightarrow\infty$. This result has two components: (i) the equilibrium is unique, and (ii) it is asymptotically stable with respect to (QLD). In doing so, we show that the convergence of Q-Learning is strongly tied to the sparsity of edges in the underlying graph.
We parameterise the network structure by the parameters of the model from which they are drawn. In the case of the Erdős-Rényi model, this is the probability $p$ of each edge. For the Stochastic Block model, this will be the probabilities $p_{c}$ of edges within each community and the probability $q$ of edges between communities. To parameterise the payoffs, we refer to the intensity of identical interests introduced in Hussain et al. (2024).
Definition 3 (Intensity of Identical Interests).
Let $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A}_{k})_{k\in\mathcal{N}},(A^{kl},A^{lk})_{(k,l)\in\mathcal{E}})$ be a network polymatrix game. Then the intensity of identical interests $\delta_{I}$ of $\mathcal{G}$ is given by
$\delta_{I}=\max_{(k,l)\in\mathcal{E}}\lVert A^{kl}+(A^{lk})^{\top}\rVert_{2}.$ (2)
Remark.
The intensity of identical interests $\delta_{I}>0$ measures the similarity between the payoffs of connected agents across all edges. A canonical example is the pairwise zero-sum game in which $A^{kl}=-(A^{lk})^{\top}$ for all $(k,l)\in\mathcal{E}$. In this case, the intensity of identical interests is zero.
To prove our main result, we make the following assumption on the network polymatrix games.
Assumption 1.
Each edge is assigned the same bimatrix game, i.e., $(A^{kl},A^{lk})=(A,B)$ for all $(k,l)\in\mathcal{E}$.
Assumption 1 is well motivated in the literature as it analyses the case where agents are engaged in the same interaction with multiple opponents, a format that is commonly used to study the emergence of cooperation (Zhang et al., 2014; Mukhopadhyay & Chakraborty, 2020) and congestion (Szabó & Szolnoki, 2015) in multi-agent systems. Indeed, the assumption of shared payoffs has led to a number of successes in studying large scale systems with many agents Perrin et al. (2020); Wu et al. (2024); Pérolat et al. (2022); Parise & Ozdaglar (2023); Hu et al. (2019). We adopt this approach, but also evaluate the validity of our claims outside of this assumption in our experiments (c.f. Figure 1).
Having specified our setting, we next determine a sufficient condition for the convergence of (QLD) in terms of the properties of the adjacency matrix $G$. All proofs are deferred to Appendix B and C.
Lemma 1.
Let $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A})_{k},(A,B)_{(k,l)\in\mathcal{E}})$ be a network polymatrix game that satisfies Assumption 1. Also let $G$ be the adjacency matrix associated with the edge-set $\mathcal{E}$. If, for each agent $k$, the exploration rate $T_{k}$ satisfies
$T_{k}>\delta_{I}\rho(G),$ (3)
the QRE $\mathbf{x}^{*}$ of the game $\mathcal{G}$ is unique. In addition, for all initial conditions, trajectories of (QLD) converge to $\mathbf{x}^{*}$.
Notice that Lemma 1 makes two claims. The first is that the QRE is unique, whilst the second shows that it is asymptotically stable under (QLD). The combination of these statements ensures that there is a unique equilibrium solution that is learned by the agents.
Discussion.
In fully connected networks, where all agents are interacting with one another, we have $\rho(G)=N-1$, which, by Lemma 1, implies that higher exploration rates are required to guarantee convergence as the number of agents increases. This agrees with previous findings in Sanders et al. (2018) and Hussain et al. (2023). However, recall from Sec. 2.2 that, as $T_{k}\rightarrow\infty$ for all agents $k$, the QRE tends towards the uniform distribution, a trivial solution in which payoffs are ignored and the game is played randomly.
From Lemma 1, it is clear that, by controlling each term in (3), (QLD) can converge even with reasonably low exploration rates [^1] The definition of ‘reasonable’ depends on the tolerable distance from the Nash Equilibrium on the part of the user. See (Hussain et al., 2024) for an upper bound on such a distance.. An example is the aforementioned pairwise zero-sum game in which $\delta_{I}=0$ and (QLD) is convergent for all $T_{k}>0$. This agrees with the result found by Leonardos et al. (2024). One may also control $\rho(G)$ by choosing networks in which the number of edges does not grow with the number of players. Such networks include lattices Szabó & Borsos (2016) or ring networks (Kleinberg et al., 2011), in which each agent $k$ is connected to agents $(k-1)\mod N$ and $(k+1)\mod N$ respectively. In this case, the agent has only two neighbours, regardless of $N$ and so the spectral radius is always $\rho(G)=2$.
In order to lift the restriction of a constant number of edges for each agent, we move towards the analysis of random networks, in which the probability of edge connections is parameterised. In doing so, we can make broader statements about the relationship between the convergence of (QLD) in many-player games and the underlying network. Our goal is to place an upper bound on the spectral radius term $\rho(G)$ in Lemma 1 in terms of the parameters of the random graph model.
Erdős-Rényi.
We begin with the Erdős-Rényi model (Erdös & Rényi, 2006), and show that the spectral radius is related to the probability of edge connections in the network. In the Erdős-Rényi model, a graph is generated by sampling each edge with a probability $p\in(0,1)$ independently. The adjacency matrix $G$ is then a random matrix with entries that are Bernoulli distributed with parameter $p$. Let the degree of a node $k$ be the number of edges $(k,l)\in\mathcal{E}$ that contain $k$. The expected degree of each node in a network drawn from the Erdős-Rényi model is therefore $p(N-1)$.
Lemma 2.
Consider the symmetric adjacency matrix $G\in\mathbb{R}^{N\times N}$ of a network drawn from the Erdős-Rényi model. Then $G$ has a zero main-diagonal and the off-diagonal entries $\{[G]_{kl}\sim\texttt{Bernoulli}(p):1\leq k<l\leq N\}$ are independent, where $p>0$. Let $\sigma=\sqrt{p(1-p)}$. For any $\epsilon>0$, the following holds almost surely as $N\to\infty$
$\rho(G)\leq(N-1)p+(2\sigma+\epsilon)\sqrt{N}.$ (C1)
The bound is dominated by the term $p(N-1)$ as $N\rightarrow\infty$, which is exactly the expected degree of a node in the Erdős-Rényi model. We now connect this result with Lemma 1 to show the convergence of (QLD).
Theorem 1.
Let $\epsilon>0$ and let $\mathcal{G}$ be a network polymatrix game satisfying Assumption 1 and suppose that the network is drawn from the Erdős-Rényi model. If
$T_{k}>\delta_{I}[(N-1)p+(2\sigma+\epsilon)\sqrt{N}],$
then it holds almost surely as $N\to\infty$ that all trajectories of (QLD) converge to a unique QRE $\mathbf{x}^{*}\in\Delta$.
Stochastic Block Model.
Networks are drawn from the Stochastic Block Model (Holland et al., 1983) by partitioning the nodes into $C$ communities. The probability of an edge between nodes $k$ and $l$ that are in the same community $c$ is given by $p_{c}\in(0,1)$, whereas the probability of an edge between nodes in different communities is given by $q\in(0,1)$. The adjacency matrix $G$ is then generated by sampling each edge independently according to Bernoulli distribution with probabilities $p_{c}$ and $q$. The expected degree of a node in community $c$ is $p_{c}(N_{c}-1)+q(N-N_{c})$, where $N_{c}$ is the number of nodes in community $c$. Notice that the Erdős-Rényi model can be seen as a limit case of the Stochastic Block Model for a single community, i.e., $N_{c}=N$. For simplicity, the following lemma assumes all communities are of equal size, specifically $N_{c}=N/C$ for all $c$. However, the proof stays valid even when the communites’ size varies.
Lemma 3.
Consider the symmetric adjacency matrix $G$ of a network drawn from the Stochastic Block Model. Then
$\begin{cases}[G]_{kl}\sim\texttt{Bernoulli}(p_{c}),&\text{ if }k\text{ and }l\in c,\\ [G]_{kl}\sim\texttt{Bernoulli}(q),&\text{ if }k\in c\text{ and }l\in c^{\prime}\neq c,\\ [G]_{kk}=0,&\text{ for }1\leq k\leq N,\end{cases}$
where the entries $\{[G]_{kl}:1\leq k<l\leq N\}$ are independent. For any $\epsilon>0$, it holds almost surely as $N\to\infty$
$\rho(G)\leq Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}+(2\sigma_{\max}+\epsilon)\sqrt{N},$ (4)
where $p_{\max}=\max\{p_{1},\ldots,p_{C}\}$ and $\sigma_{\max}=\max\{\sqrt{p_{1}(1-p_{1})},\ldots,\sqrt{p_{C}(1-p_{C})},\sqrt{q(1-q)}\}$.
Remark.
The generalizability of Lemmas 2 and 3 warrants clarification. First, the $\epsilon\sqrt{N}$ term can be replaced by $\epsilon N^{1/4+\delta}\ln N$ for any $\delta>0$, as can be seen in Appendix C.
By considering the $O(N)$ dominant term $Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}$ in the bound (4), we notice that it can be rearranged to $p_{\max}(\frac{N}{C}-1)+q(N-\frac{N}{C})$, which is exactly the expected degree of a node in a Stochastic Block Model with $p_{c}=p_{\max}$ and $N_{c}=N/C$ for all subcommunities $c$. We apply this in the following theorem to connect the probability of edge connections in the graph with the convergence of (QLD).
Theorem 2.
Let $\epsilon>0$ and let $\mathcal{G}$ be a network polymatrix game satisfying Assumption 1. Suppose that the network is drawn from the Stochastic Block model. Finally, let $p_{\max},\sigma_{\max}$ be defined as in Lemma 3. If
$T_{k}>\delta_{I}\left[Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}+(2\sigma_{\max}+\epsilon)\sqrt{N}\right],$
the trajectories of (QLD) converge to a unique QRE $\mathbf{x}^{*}\in\Delta$ almost surely as $N\to\infty$.
We conclude our results with a point of independent interest, in which we prove sufficient conditions on the uniqueness of the Nash Equilibrium in games played on networks drawn from the Erdős-Rényi and Stochastic Block models. These hold as corollaries of Theorems 1 and 2 and by recognising that the Nash Equilibrium and QRE coincide in the limit $T_{k}\rightarrow 0$ for all $k$.
Corollary 1.
Let $\epsilon>0$ and $\mathcal{G}$ be a network polymatrix game satisfying Assumption 1 and $G$ the adjacency matrix of the network. The Nash Equilibrium is unique almost surely as $N\to\infty$ if the network is drawn from the (i) Erdős-Rényi model and
$0>\delta_{I}\left[(N-1)p+(2\sigma+\epsilon)\sqrt{N}\right],$
where $\sigma,p$ are defined as in Lemma 2 or (ii) the Stochastic Block model and
$0>\delta_{I}\left[Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}+(2\sigma_{\max}+\epsilon)\sqrt{N}\right],$
where $p_{\max},q,\sigma_{\max}$ are defined as in Lemma 3.
4 Convergence of Q-Learning in Random Graphs: experimental evaluation
Figure 1: Proportion of converged runs after simulating (QLD) for choices of $(p,T)$ in network games where the network is drawn from the Erdős-Rényi model. We depict the (top) Network Sato Game (middle) Network Shapley Game with $\beta=0.2$ (bottom) Conflict Network Game, each with varying $N$ across heatmaps. For the Network Sato Game ($\delta_{I}=0.2$), we also depict the lower bound predicted by Theorem 1 whilst for the Network Shapley Game $\delta_{I}=2.0$, the bound exceeds the range of the plot and, due to the random nature of the payoffs, it is not possible to depict the bound for the Conflict Network Game.
Figure 2: Variation of the empirical convergence boundary in the (Left) Network Sato Game (Right) Network Shapley Game as $N$ increases, for different values of $p$. The $y$-axis shows the smallest exploration rate for which all simulations converged. In the Network Sato Game, we also plot the theoretical bound as a dashed line.
Figure 3: Proportion of converged runs after simulating (QLD) for choices of $(p,T)$ in a Network Sato Game where the network is drawn from the Stochastic Block Model for varying choices of number of agents $N$ and $q$. The theoretical lower bound from Theorem 2 is plotted in white.
The aim of this section is to understand the implications of Theorems 1 and 2 in games with a finite number of agents. To do this, we simulate the Q-Learning algorithm described in Sec. 2.3 on network games, where the network is drawn from the Erdős-Rényi or Stochastic Block models. We show that, as the sparsity of the network decreases, higher exploration rates are required to guarantee the convergence of Q-Learning, by finding a boundary between stable and unstable dynamics. We analyze how our lower bound and the true boundary evolve with $N$ and provide empirical evidence that our results may extend beyond the assumptions used to derive the theoretical bounds.
Remark.
In our experiments, we assume all agents have the same exploration rate $T$ and thus omit the $k$ notation. As our results are lower bounds across all $k$, this is without loss of generality.
Game Models.
We first simulate Q-Learning in the Network Shapley and Network Sato Games, which are extensions of classical bimatrix games, introduced in Shapley (2016) and Sato et al. (2002) respectively, to the network polymatrix setting. In both cases, each edge defines the same bimatrix game, $(A^{kl},A^{lk})=(A,B)\;\forall(k,l)\in\mathcal{E}$ where in the Network Shapley Game:
$A=\begin{pmatrix}1&0&\beta\\ \beta&1&0\\ 0&\beta&1\end{pmatrix},\,B=\begin{pmatrix}-\beta&1&0\\ 0&-\beta&1\\ 1&0&-\beta\end{pmatrix},$
for $\beta\in(0,1)$; and in the Network Sato Game:
$A=\begin{pmatrix}\epsilon_{X}&-1&1\\ 1&\epsilon_{X}&-1\\ -1&1&\epsilon_{X}\end{pmatrix},\,B=\begin{pmatrix}\epsilon_{Y}&-1&1\\ 1&\epsilon_{Y}&-1\\ -1&1&\epsilon_{Y}\end{pmatrix},$
where $\epsilon_{X},\epsilon_{Y}\in\mathbb{R}$. In our experiments, we fix $\epsilon_{X}=0.5,\epsilon_{Y}=-0.3$ and $\beta=0.2$.
Previous works have shown that orbits of (QLD) in these games can exhibit complex behaviours such as limit cycles and chaos for low exploration rates.
Erdős-Rényi Model.
In Figure 1, we examine the convergence of Q-Learning in the case that the network is drawn from the Erdős-Rényi model. We vary the probability $p$ of edge connections and the exploration rate $T$ for different numbers $N$ of agents. For each parameter choice, we simulate Q-Learning for 3000 rounds and for 50 initial conditions and record whether the dynamics converged to a fixed point. We determine convergence by the variance of the agent strategies in the final 300 rounds of the simulation.
In the case of the Network Sato Game, we also depict the lower bound predicted by Theorem 1. We observe that the theoretical bound accurately models the increase in the exploration rate required for convergence as $p$ increases, except a small region as $p\rightarrow 1$. This is due to the $\sigma$ term in the bound, which vanishes as $N\rightarrow\infty$ and so the bound becomes dominated by the $p(N-1)$ term. We also observe that, as $N$ increases, the region of non-convergence expands rapidly for high $p$, whilst remaining relatively fixed for low $p$. This is consistent with previous results in the literature (Hussain et al., 2024; Parise & Ozdaglar, 2019). We expand on this in Figure 2 in which we plot the boundary as $N$ increases for different values of $p$. We observe that the rate at which the boundary grows increases with $p$.
Figure 1 also shows the influence of the $\delta_{I}$ term in the comparison of the Network Sato ($\delta_{I}=0.2$) and Network Shapley ($\delta_{I}=2.0$) Games. In particular, the boundary, for any given parameter choice, occurs at higher exploration rates in the Network Shapley Game.
Finally, we consider the Conflict Network Game (Ewerhart & Valkanova, 2020), where the payoff matrices need not be the same along each edge. A network polymatrix game is a conflict network game if each bimatrix game $(A^{kl},A^{lk})$ satisfies $(A^{kl})_{ij}=v^{k}(P^{kl})_{ij}-c^{kl}_{i}$ and $(A^{lk})_{ji}=v^{j}(P^{lk})_{ji}-c^{lk}_{j}$, where $v_{k},v_{l}>0$, $c_{kl}\in\mathbb{R}^{n_{k}},c_{lk}\in\mathbb{R}^{n_{l}}$ and $(P^{kl})_{ij}+(P^{lk})_{ji}=1$ for all $i\in S_{k},j\in S_{l}$. As pointed out in Ewerhart & Valkanova (2020), if $v_{k}=v_{l}$ for all $(k,l)\in\mathcal{E}$, the conflict network is equivalent to a network zero sum game. To generate Figure 1, we independently sample each $v_{k}$ from the uniform distribution over $[0,1]$. We then randomly generate The matrix $P^{kl}$ and subsequently generate $(A^{kl},A^{lk})$ such that the conflict network condition is satisfied. We find that, although the boundary is not as clearly defined due to the random nature of the payoffs, the implication that convergence occurs with lower exploration rates in sparser networks still holds.
Stochastic Block Model.
We repeat our experiments for the case that the network is drawn from a Stochastic Block Model. In Figure 3, we depict the proportion of converged experiments in the Network Sato Game, with Network Shapley depicted in the Appendix (Figure 4). Due to the large number of parameters in the Stochastic Block Model, we fix the number of agents in each community to $N_{c}=5$ and suppose that $p_{c}=p$ for all communities $c$. We then determine the convergence of (QLD) for different values of $p$ and $T$ and vary across plots $q$ and $N$. We observe that whilst each of these parameters plays a role in determining asymptotic behaviour, it is clear that convergence occurs less frequently in denser graphs, i.e., as $p$ and $q$ increase.
5 Conclusion
This paper examined Q-Learning dynamics in network polymatrix games through the lens of random networks. In doing so, we showed an explicit relationship between the asymptotic convergence of Q-Learning and the sparsity of the underlying network. We provided theoretical lower bounds on the exploration rates required for convergence in network polymatrix games where the network is drawn from the Erdős-Rényi or Stochastic Block models and showed that the convergence may be guaranteed in games with many agents, so long as the sparsity of the network is carefully controlled.
Future Work.
In our numerical simulations, we observed that the implications of our bounds may hold in certain cases outside of the assumptions used to derive them. Future work should therefore examine whether our results can hold in various classes of network games, such as games with continuous action sets or games with more complex payoff structures. In addition, whilst our work considers the repeated play of a matrix game, future work should consider our results with the introduction of a state variable, as in the case of Markov Games. Doing so opens up the possibility for strong guarantees to be placed in large multi-agent systems which may be deployed in real world settings.
Impact Statement
Our work aims to advance understanding into the foundations of multi-agent reinforcement learning. As our results are foundational, there are no specific societal or ethical consequences of our work that must be highlighted here.
Appendix A Variational Inequalities and Monotone Games
The main idea of our convergence proof is to show that, under the conditions (3), the corresponding network game is a strictly monotone game (Melo, 2018; Parise & Ozdaglar, 2019; Hadikhanloo et al., 2022; Sorin & Wan, 2016). In such games, it is known that the equilibrium solution is unique (Melo, 2021; Facchinei & Pang, 2004). It is further known, from (Hussain et al., 2023), that (QLD) converges asymptotically in monotone games. We leverage this result to prove Lemma 1.
We begin by framing game theoretic concepts in the language of variational inequalities.
Definition 4 (Variational Inequality).
Consider a set $\mathcal{X}\subset\mathbb{R}^{d}$ and a map $F\,:\mathcal{X}\rightarrow\mathbb{R}^{d}$. The Variational Inequality (VI) problem $VI(\mathcal{X},F)$ is given as
[…]
We say that $\mathbf{x}^{*}\in\mathcal{X}$ belongs to the set of solutions to a variational inequality problem $VI(\mathcal{X},F)$ if it satisfies (5).
We now wish to reformulate the problem of finding Quantal Response Equilibria, or Nash Equilibria, as a problem of solving a variational inequality of a particular form. In such a case, the set $\mathcal{X}$ is identified with the joint simplex $\Delta$. The map $F$ is identified with the pseudo-gradient map of the game.
Definition 5 (Pseudo-Gradient Map).
The $\mathcal{G}$ be a network polymatrix game with payoff functions $(u_{k})_{k\in\mathcal{N}}$. Then, the pseudo-gradient map of $\mathcal{G}$ is $F:\mathbf{x}\mapsto(-D_{\mathbf{x}_{k}}u_{k}(\mathbf{x}_{k},\mathbf{x}_{-k}))_{k\in\mathcal{N}}$.
This reformulation has been used, for example, by (Melo, 2021) to show that a QRE of a game can be found by solving a variational inequality of a particular form. We reformulate their theorem for the particular case of network polymatrix games.
Lemma 4 (Melo (2021)).
Consider a game $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A})_{k\in\mathcal{N}},(A^{kl},A^{lk})_{(k,l)\in\mathcal{E}})$ and for any $T_{1},\ldots,T_{N}>0$, let the regularised game $\mathcal{G}^{H}$ be the network game in which the payoff $u_{k}^{H}$ to each agent $k$ is given by
$u_{k}^{H}(\mathbf{x}_{k},\mathbf{x}_{-k})=\sum_{l:(k,l)\in\mathcal{E}}\mathbf{x}_{k}^{\top}A^{kl}\mathbf{x}_{l}-T_{k}\langle\mathbf{x}_{k},\ln\mathbf{x}_{k}\rangle.$ (6)
Now, let $F^{H}$ be the pseudo-gradient map of $\mathcal{G}^{H}$. Then $\mathbf{x}^{*}\in\Delta$ is a QRE of $\mathcal{G}$ if and only if $\mathbf{x}^{*}$ is a solution to $VI(\Delta,F^{H})$. .
In fact, this same lemma can be used to show that $\mathbf{x}^{*}$ is a QRE of $\mathcal{G}$ if and only if it is a Nash Equilibrium of $\mathcal{G}^{H}$. This concept of regularised games has also been used to show connections between the replicator dynamics (Maynard Smith, 1974) and Q-Learning dynamics in (Leonardos & Piliouras, 2022), and to design algorithms for Nash Equilibrium seeking (Gemp et al., 2022). We require one final component from the study of variational inequalities which is often used to study uniqueness of equilibrium solutions: monotonicity (Parise & Ozdaglar, 2019; Melo, 2018).
Definition 6 (Monotone Games).
Let $\mathcal{G}$ be a network polymatrix game with pseudo-gradient map $F$. $\mathcal{G}$ is said to be
  1. 1.
    Monotone if, for all $\mathbf{x},\mathbf{y}\in\Delta$,
    $\langle F(\mathbf{x})-F(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle\geq 0.$
  2. 2.
    Strictly Monotone if, for all $\mathbf{x},\mathbf{y}\in\Delta$,
    $\langle F(\mathbf{x})-F(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle>0.$
  3. 3.
    Strongly Monotone with constant $\alpha>0$ if, for all $\mathbf{x},\mathbf{y}\in\Delta$,
    $\langle F(\mathbf{x})-F(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle\geq\alpha||\mathbf{x}-\mathbf{y}||^{2}_{2}.$
By proving monotonicity properties of a game, we can leverage the following results from literature.
Lemma 5 (Melo (2021)).
Consider a game $\mathcal{G}$ and for any $T_{1},\ldots,T_{N}>0$, let $F$ be the pseudo-gradient map of $\mathcal{G}^{H}$. $\mathcal{G}$ has a unique QRE $\mathbf{x}^{*}\in\Delta$ if $\mathcal{G}^{H}$ is strongly monotone with any $\alpha>0$.
Lemma 6 (Hussain et al. (2023)).
If the game $G$ is monotone, then the Q-Learning Dynamics (QLD) converges to a unique QRE $\mathbf{x}^{*}\in\Delta$ with any positive exploration rates $T_{1},\ldots,T_{N}>0$.
Finally, note that a map $g:\Delta\rightarrow\mathbb{R}$ is strongly convex with constant $\alpha$ if, for all $\mathbf{x},\mathbf{y}\in\Delta$
$g(\mathbf{y})\geq g(\mathbf{x})+Dg(\mathbf{x})^{\top}(\mathbf{y}-\mathbf{x})+\frac{\alpha}{2}\lVert\mathbf{x}-\mathbf{y}\rVert^{2}_{2}.$
If the map $g(\mathbf{x})$ is twice-differentiable, then it is strongly convex if its Hessian $D^{2}_{\mathbf{x}}g(\mathbf{x})$ is strongly positive definite with constant $\alpha$. Thus, all eigenvalues of $D^{2}_{\mathbf{x}}g(\mathbf{x})$ are larger than $\alpha$. It also holds that, if $D^{2}_{\mathbf{x}}g(\mathbf{x})$ is strongly positive definite, the gradient $D_{\mathbf{x}}g(\mathbf{x})$ is strongly monotone. To apply this in our setting, we use the following result.
Proposition 1 (Melo (2021)).
The function $g(\mathbf{x}_{k})=T_{k}\langle\mathbf{x}_{k},\ln\mathbf{x}_{k}\rangle$ is strongly convex with constant $T_{k}$.
A.1 Matrix Norms
We close with a few properties of matrices that are useful towards our decomposition of the payoff matrices and the graph adjacency matrix.
Proposition 2.
For any matrix $A$, $\lVert A\rVert_{2}=\sqrt{\lambda_{\max}(A^{\top}A)}$, where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue of a matrix. If, in addition, the matrix is symmetric, $\lVert A\rVert_{2}=\rho(A)$, where $\rho(A)$ is the spectral radius of $A$.
Proposition 3 (Weyl’s Inequality).
Let $J=D+N$ where $D$ and $N$ are symmetric matrices. Then it holds that
$\lambda_{\min}(J)\geq\lambda_{\min}(D)+\lambda_{\min}(N),$
where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue of a matrix.
Proposition 4.
Let $G,A$ be matrices and let $\otimes$ denote the Kronecker product. Then
$\lVert G\otimes A\rVert_{2}=\lVert G\rVert_{2}\lVert A\rVert_{2}$ (7)
Proposition 5.
Let $A$ be a symmetric matrix. Then
$|\lambda_{\min}(A)|\leq\rho(A)$
Lemma 7.
[…]
$\lVert A\rVert_{2}\leq\lVert B\rVert_{2}.$
Proof.
Let +$\mathbb{R}^{n}_{+}$ be the set of $n$-dimensional nonnegative vectors, i.e. with nonnegative entries. As $A$ and $B$ have nonnegative entries, we deduce that
$\lVert A\rVert_{2}=\sup_{\mathbf{x}\in\mathbb{R}^{n}}\lVert A\mathbf{x}\rVert_{2}=\sup_{\mathbf{x}\in\mathbb{R}^{n}_{+}}\lVert A\mathbf{x}\rVert_{2},$
$\lVert B\rVert_{2}=\sup_{\mathbf{x}\in\mathbb{R}^{n}}\lVert B\mathbf{x}\rVert_{2}=\sup_{\mathbf{x}\in\mathbb{R}^{n}_{+}}\lVert B\mathbf{x}\rVert_{2}.$
Further, for any $\mathbf{x}\in\mathbb{R}^{n}_{+}$, it holds that
$\lVert B\mathbf{x}\rVert_{2}=\lVert(B-A)\mathbf{x}+A\mathbf{x}\rVert_{2}\geq\lVert A\mathbf{x}\rVert_{2},$
as $(B-A)\mathbf{x}$ is nonegative. This increases the norm of any nonnegative vector $A\mathbf{x}$. Taking supremum in the above equation, we obtain that
$\sup_{\mathbf{x}\in\mathbb{R}^{n}}\lVert A\mathbf{x}\rVert_{2}=\sup_{\mathbf{x}\in\mathbb{R}_{+}^{n}}\lVert A\mathbf{x}\rVert_{2}\leq\sup_{\mathbf{x}\in\mathbb{R}_{+}^{n}}\lVert B\mathbf{x}\rVert_{2}=\sup_{\mathbf{x}\in\mathbb{R}^{n}}\lVert B\mathbf{x}\rVert_{2}.$
Appendix B Proof of Lemma 1
The proof takes the following steps. We first decompose the derivative of the pseudo-gradient, which we dub the pseudo-jacobian, of the regularised game $\mathcal{G}^{H}$ into a term involving payoffs and a term involving exploration rates. In doing so, we can determine how the exploration rates should be balanced so that the pseudo-jacobian is positive definite, which yields monotonicity of the game. We then further decompose the payoff term into terms involving the payoff matrices and the network adjacency matrix. This exposes the connection between each of the three quantities: exploration rate, payoff matrices and network connectivity.
Proof of Lemma 1.
Let $F$ be the pseudo-gradient of the regularised game $\mathcal{G}^{H}$. We define the pseudo-jacobian as the derivative of $F$, given by
$[J(\mathbf{x})]_{k,l}=D_{\mathbf{x}_{l}}F_{k}(\mathbf{x})$
It holds that if $\frac{J(\mathbf{x})+J^{\top}(\mathbf{x})}{2}$ is positive definite for all $\mathbf{x}\in\Delta$ then $F(\mathbf{x})$ is monotone. We decompose $J$ as
$J(\mathbf{x})=D(\mathbf{x})+N(\mathbf{x}),$
where $D(\mathbf{x})$ is a block diagonal matrix with $-D^{2}_{\mathbf{x}_{k}\mathbf{x}_{k}}u_{k}^{H}(\mathbf{x}_{k},\mathbf{x}_{-k})$ along the diagonal. $N(\mathbf{x})$ is an off-diagonal block matrix with
$[N(\mathbf{x})]_{k,l}=\begin{cases}-D_{\mathbf{x}_{k},\mathbf{x}_{l}}u_{k}^{H}(\mathbf{x}_{k},\mathbf{x}_{-k})&\text{ if }(k,l)\in\mathcal{E}\\ \mathbf{0}&\text{ otherwise}\end{cases}.$
Now notice that $-u_{k}^{H}(\mathbf{x}_{k},\mathbf{x}_{-k})=T_{k}\langle\mathbf{x}_{k},\ln\mathbf{x}_{-k}\rangle-\sum_{l:(k,l)\in\mathcal{E}}\mathbf{x}_{k}^{\top}A^{kl}\mathbf{x}_{l}$. Therefore, $D(x)$ is simply the Hession of the entropy regularisation term $T_{k}\langle\mathbf{x}_{k},\ln\mathbf{x}_{k}\rangle$. From Proposition 1, it holds then that $D(\mathbf{x})$ is strongly positive definite with constant $T=\min_{k}T_{k}$. Let $\bar{J}(\mathbf{x})$ be defined as
$\bar{J}(\mathbf{x})=D(\mathbf{x})+\frac{N(\mathbf{x})+N^{\top}(\mathbf{x})}{2}.$
In words, $\bar{J}(\mathbf{x})$ is the symmetric component of $J(\mathbf{x})$. We may now use Weyl’s inequality to write
$\lambda_{\min}(\bar{J}(\mathbf{x}))$ $\geq\lambda_{\min}(D(\mathbf{x}))+\lambda_{\min}\left(\frac{N(\mathbf{x})+N(\mathbf{x})^{\top}}{2}\right)$
$\geq T-\rho\left(\frac{N(\mathbf{x})+N(\mathbf{x})^{\top}}{2}\right)$
$=T-\frac{1}{2}\lVert N(\mathbf{x})+N(\mathbf{x})^{\top}\rVert_{2}$
To determine $\lVert N(\mathbf{x})+N(\mathbf{x})^{\top}\rVert_{2}$, we first notice that, in network polymatrix games each block of $N(\mathbf{x})$ is given by
$[N(\mathbf{x})]_{k,l}=\begin{cases}-A^{kl}&\text{ if }(k,l)\in\mathcal{E}\\ \mathbf{0}\text{ otherwise.}\end{cases}$
whilst
$[N(\mathbf{x})^{\top}]_{k,l}=\begin{cases}-(A^{lk})^{\top}&\text{ if }(k,l)\in\mathcal{E}\\ \mathbf{0}\text{ otherwise.}\end{cases}$
To write this in the form of a kronecker product, we leverage Assumption 1, namely that each edge corresponds to the same bimatrix game with payoff matrices $(A,B)$. We decompose each edge into a half-edge along which $A$ is played, and a half-edge along which $B$ is played. In doing so, we may decompose the adjacency matrix $G=G_{k\rightarrow l}+G_{l\rightarrow k}$. The non-zero elements of $G_{k\rightarrow l}$ correspond to half edges along which $A$ is played, $G_{l\rightarrow k}$ denote the half-edges along which $B$ is played. With this definition in place, we may write
$N(\mathbf{x})+N(\mathbf{x})^{\top}=-(A+B^{\top})\otimes G_{k\rightarrow l}-(A^{\top}+B)\otimes G_{l\rightarrow k}$
Then, by Proposition 4
$\frac{1}{2}\lVert N(\mathbf{x})+N(\mathbf{x})^{\top}\rVert_{2}$ $=\frac{1}{2}\lVert(A+B^{\top})\otimes G_{k\rightarrow l}+(A^{\top}+B)\otimes G_{l\rightarrow k}\rVert_{2}$
$\leq\frac{1}{2}\lVert(A+B^{\top})\otimes G_{k\rightarrow l}\rVert_{2}+\frac{1}{2}\lVert(A^{\top}+B)\otimes G_{l\rightarrow k}\rVert_{2}$
$=\frac{1}{2}\lVert A+B^{\top}\rVert_{2}(\lVert G_{k\rightarrow l}\rVert_{2}+\lVert G_{l\rightarrow k}\rVert_{2})$
$\leq\lVert A+B^{\top}\rVert_{2}\lVert G\rVert_{2}$
$=\delta_{I}\rho(G)$
where in the final inequality, we use Lemma 7. From this we may establish that
$\lambda_{\min}(\bar{J}(x))$ $\geq T-\delta_{I}\rho(G)$
Therefore, if $T>\delta_{I}\rho_{G}$, the pseudo-jacobian is strongly positive definite with constant $T-\delta_{I}\rho(G)$, from which it is established that $\mathcal{G}^{H}$ is strongly monotone with the same constant. From Lemma 5 (Melo, 2021) the QRE is unique, and from Lemma 6, it is asymptotically stable under (QLD). ∎
Appendix C Proofs of Lemmas 2 and 3
We now focus on establishing upper bounds on the spectral radius when the network is drawn from the Erdős-Rényi or Stochastic Block models. The proof idea is to decompose $G$ into $\mathbb{E}[G]$ and $\tilde{G}=G-\mathbb{E}[G]$. Then, $\rho(\mathbb{E}[G])$ is deterministic and can be computed with techniques from linear algebra. For $\rho(\tilde{G})$, we require the next result, which is a particular case of Theorem 1.4 from Vu (2007). Notation-wise, let $I_{N}$ be the identity matrix and $J_{N}$ be the matrix whose elements are all ones.
Lemma 8.
[…]
$\rho(A)\leq 2\sigma_{\max}\sqrt{N}+C\sqrt{K\sigma_{\max}}N^{1/4}\ln N.$
The constant $C$ can be removed. In particular, for any $\epsilon>0$,
$\rho(A)\leq(2\sigma_{\max}+\epsilon)\sqrt{N}\,\text{ almost surely as }N\to\infty.$
Proof of Lemma 2.
For the deterministic part, we have that
$\mathbb{E}[G]=\begin{pmatrix}0&p&p&\cdots&p\\ p&0&p&\cdots&p\\ p&p&0&\cdots&p\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ p&p&p&\cdots&0\end{pmatrix}=pJ_{N}-pI_{N},$ (8)
with eigenvalues $\lambda_{1}=(N-1)p$ and $\lambda_{2}=\ldots=\lambda_{N}=-p$. Hence $\rho(\mathbb{E}[G])=(N-1)p$. Further, by applying Lemma 8, we obtain that almost surely as $N\to\infty$
$\rho\left(\tilde{G}\right)\leq\left(2\sqrt{p(1-p)}+\epsilon\right)\sqrt{N}.$
The result follows from noting that
$\rho(G)\leq\rho(\mathbb{E}[G])+\rho\left(\tilde{G}\right)=(N-1)p+\left(2\sqrt{p(1-p)}+\epsilon\right)\sqrt{N}$
holds almost surely as $N\to\infty$. ∎
Proof of Lemma 3.
For the deterministic part, we have that
$\mathbb{E}[G]=\begin{pmatrix}P_{1}&Q&Q&\cdots&Q\\ Q&P_{2}&Q&\cdots&Q\\ Q&Q&P_{3}&\cdots&Q\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ Q&Q&Q&\cdots&P_{C}\end{pmatrix},$
where $Q=qJ_{N}$ and $P_{l}=pJ_{m}-pI_{m}$ for $1\leq l\leq C$, as in (8). Thus
$\rho\left(\mathbb{E}[G]\right)\leq\lVert\mathbb{E}[G]\rVert_{1}=$ $(N-N/C)q+\max_{1\leq l\leq C}{\{(N/C-1)p_{l}\}}=(N-N/C)q+(N/C-1)p_{\text{max}}$
$=$ $Nq+N(p_{\text{max}}-q)/C-p_{\text{max}},$
where $p_{\text{max}}=\max_{1\leq l\leq C}p_{l}$. For $\tilde{G}$, let $\sigma_{\text{max}}=\max{\{\sqrt{p_{1}(1-p_{1})},\ldots,\allowbreak\sqrt{p_{C}(1-p_{C})},\sqrt{q(1-q)}\}}$. By applying Lemma 8, we obtain that almost surely as $N\to\infty$
$\rho\left(\tilde{G}\right)\leq\left(2\sigma_{\text{max}}+\epsilon\right)\sqrt{N}.$
The result follows by noting that $\rho(G)\leq\rho(\mathbb{E}[G])+\rho\left(\tilde{G}\right)$ and combining the two upper bounds above. ∎
Figure 4: Proportion of converged runs after simulating (QLD) for choices of $(p,T)$ in a Network Shapley Game where the network is drawn from the Stochastic Block Model for varying choices of number of agents $N$ and $q$. We find that each parameter both the random graph model parameters $p,q,N$ and the exploration rate $T$ influence the stability of (QLD).